# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for MLLM Hallucination Reduction

### v3

@@ 4.3 Ablation Analyses on Component
Algorithm Data Size Feedback Reward-Free Hallucination Mitigation Causal Alignment
LLaVA-v1.5 (Liu et al., 2023b) - -
+ RLHF (Sun et al., 2024) 122k self-reward
+ RLAIF (Yu et al., 2025) 16k LLaVA-Next
+ HALVA (Sarkar et al., 2025a) 22k GPT-4V
+ DPO (Li et al., 2024) 5k self-reward
+ CHiP-DPO (Fu et al., 2025) 5k self-reward
+ OPA-DPO (Yang et al., 2025) 4.8k GPT-4V
@@
4.4 Ablation Analyses on Preference Scale Impact
We investigate how the scale of preference data affects hallucination suppression in TARS by training on subsets of the 4.8k dataset and comparing with standard DPO (Figure 5). TARS consistently achieves lower hallucination rates across all scales and shows sharper improvements in early stages. From 0 to 1.8k examples, the 7B and 13B variants reduce hallucinations by over 18 and 15 percentage points, respectively. While the gains taper beyond 3.6k, performance remains stable, indicating strong data efficiency and effective use of limited supervision compared to DPO.
4.5 Stability of Semantic Representations
We analyze how preference optimization reshapes hidden-state distributions in Figure 6.
(1) Disentanglement of hallucinated and preference characteristics. TARS yields a more structured latent space, where hallucinated and preference-aligned representations are separated. This separation indicates that hallucinations are not artifacts of spurious preference associations. In contrast, DPO exhibits entangled clusters, where hallucinated points are interwoven with preference representations, suggesting that DPO-trained models overfit superficial signals.
(2) Selective alignment with non-hallucinated features. TARS selectively aligns non-hallucinated responses with preference features while isolating hallucinated content in the representation space. This alignment distinguishes TARS from both the scattered representations in base LLaVA and the feature entanglement in DPO models. Our findings show TARS creates a semantically faithful space by reinforcing only factually grounded responses that match learned preferences, avoiding amplification of spurious preference correlations.
5 Related Work
@@
Multimodal large language models (MLLMs) extend LLMs by integrating visual inputs to support multimodal reasoning (Chen et al., 2024b; Feng et al., 2025; Jain et al., 2024). Typically, visual features are extracted by a vision encoder, aligned through a connector, and processed by the LLM (Liu et al., 2023b; Parekh et al., 2024). Despite strong performance, MLLMs often produce factually incorrect or visually ungrounded outputs, undermining reliability (Bai et al., 2024; Chen et al., 2025). This issue is more severe than in unimodal LLMs (Chen et al., 2024d; Jiang et al., 2024a), mainly due to modality imbalance (Ma et al., 2024; He et al., 2024) and ineffective fusion (Bellagente et al., 2023; Ji et al., 2023). Recent studies attribute these failures to persistent misalignment between multimodal representations and human expectations, rather than model capacity (Chen et al., 2024c; Liu et al., 2024c; Ruan et al., 2025).
@@
A key bottleneck in addressing MLLM hallucinations lies in aligning model outputs with human preferences for factual consistency. Unlike knowledge-intensive pretraining (Chang et al., 2024; McKinzie et al., 2024) and instruction tuning (Chen et al., 2024a; Liu et al., 2023b), recent methods typically leverage small-scale human preference data refined via reinforcement learning (Yu et al., 2024; Casper et al., 2023). Direct preference optimization (DPO)(Rafailov et al., 2023; Pi et al., 2024) has become a leading approach due to its simplicity and effectiveness, demonstrated in ChiP-DPO(Fu et al., 2025) and OPA-DPO (Yang et al., 2025). However, DPO’s reliance on limited data can cause overfitting to superficial linguistic cues (Setlur et al., 2024; Fu et al., 2025), leading to distributional rigidity and reduced adaptability to modality-specific semantics (Ouali et al., 2024; Song et al., 2024). These limitations call for more adaptive alignment strategies that capture token-level variability and cross-modal dependencies.
@@
To address these challenges, we propose a token-adaptive min-max alignment strategy that enhances preference learning without relying on high-resource expert feedback (e.g., GPT-4V). Using only a small public preference dataset, our method effectively mitigates hallucinations and consistently outperforms RL-based baselines across benchmarks. Table 3 compares preference optimization methods in terms of data scale, supervision, and alignment performance.
@@ 6 Conclusion
In this work, we introduce TARS, a novel lightweight strategy that reformulates direct preference optimization (DPO) as a min-max objective. TARS maximizes token-level distributional shifts under semantic constraints to simulate alignment uncertainty, while simultaneously minimizing the expected preference loss under these controlled perturbations. This formulation encourages the model to align more faithfully with causally grounded visual cues rather than overfit to superficial textual correlations, effectively mitigating hallucinations. TARS achieves strong hallucination suppression and consistently outperforms prior methods across most standard evaluation metrics, despite using only 4.8k public preference samples and no expert-labeled feedback or large-scale teacher models. Empirical results underscore the effectiveness of token-level alignment strategies for mitigating hallucinations in low-supervision settings.
@@
  • Section A details model configurations, training settings for DPO and TARS, and token perturbation procedures.
  • Section B presents the min-max optimization algorithm of TARS in pseudocode form.
  • Section C includes extended ablations analyzing the perturbation magnitude and spectral regularization strength.
  • Section D reports additional benchmark results and fine-grained hallucination metrics.
  • Section E discusses model behavior, including sensitivity and design insights.
  • Section F showcases qualitative comparisons on representative examples.
Appendix A Implementation Details
A.1 Base Model Setups
We evaluate our method on LLaVA-v1.5 (Liu et al., 2023b) models with 7B and 13B parameters. LLaVA-v1.5 adopts Vicuna-7B/13B (Chiang et al., 2023) as the language backbone and CLIP-ViT-L/14  (Radford et al., 2021) as the vision encoder. The vision encoder also serves as the similarity function $\mathcal{G}(\cdot)$ used in Eq. (9) to compute alignment between visual inputs and text tokens. All experiments are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
A.2 DPO Training Setups
@@
4:     Compute token-level visual relevance:
$P_{i}=\mathcal{G}_{v}(x)\cdot\mathcal{G}_{t}(q_{i})^{T}.$ (13)
5:     Estimate model confidence margin:
$\Delta P=\max_{j}P_{j}-\max_{k\neq j}P_{k}.$ (14)
6:     Determine adaptive perturbation budget:
$N_{t}=\left\lfloor\omega\cdot\Delta P^{-1}\right\rfloor+1.$ (15)
7:     Select visually agnostic tokens:
$\mathcal{A}=\text{Top}_{N_{t}}(-P).$ (16)
8:     Apply controlled perturbation to obtain $\varphi(q)$:
$\varphi(q)=\left\{\mathbb{I}[i\in\mathcal{A}(x,q)]\cdot\varphi(q_{i})+\mathbb{I}[i\notin\mathcal{A}(x,q)]\cdot q_{i}\right\}_{i=1}^{|q|}.$ (17)
@@
10:     Compute the preference alignment loss via DPO:
$\mathcal{L}_{\text{DPO}}=-\log\sigma\left(\alpha\log\frac{\pi_{\theta}(y_{w}\mid x,\varphi(q))}{\pi_{\text{ref}}(y_{w}\mid x,\varphi(q))}-\alpha\log\frac{\pi_{\theta}(y_{r}\mid x,\varphi(q))}{\pi_{\text{ref}}(y_{r}\mid x,\varphi(q))}\right).$ (18)
11:     Apply frequency-domain regularization:
$\mathcal{L}_{\text{freq}}=-\log\sigma\left(\beta\cdot\left[\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{w})\right)}-\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{r})\right)}\right]\right).$ (19)
12:     Compute final objective:
$\mathcal{L}_{\text{TARS}}=\mathcal{L}_{\text{DPO}}+\lambda\cdot\mathcal{L}_{\text{freq}}.$ (20)
@@
A.3 Evaluation Benchmark Setups
We follow the original evaluation settings and benchmark splits for AMBER (Wang et al., 2023), MMHal (Sun et al., 2023), and OBJHal (Yu et al., 2024) as specified in their respective papers. For POPE (Li et al., 2023), we construct a new benchmark of 9,000 VQA pairs by sampling using the popular, random, and adversarial strategies.
For evaluation metrics, we adopt four response-level hallucination measures across different benchmarks: CHAIR (Rohrbach et al., 2018) for object hallucination detection, object coverage (Cover) for completeness measurement, response-level hallucination rate (Hal-Rate) for overall hallucination assessment, sentence-level hallucination rate (CRs) for holistic response evaluation, and object mention-level hallucination rate (CRi) for fine-grained object-level analysis.
For evaluation feedback collection, we employ the en-core-web-lg English NLP pipeline for AMBER to extract structured semantic cues as lightweight and reproducible evaluators. For MMHal and OBJHal, we utilize the expert GPT-4V model (Hurst et al., 2024) (gpt-4-1106-vision-preview) for feedback evaluation, following the established protocols of benchmarks.
Appendix B Algorithm Flowchart
Table 4: Ablation study on the effect of token-level perturbation magnitude $\omega$ in TARS. We evaluate how varying the perturbation ratio influences hallucination suppression, semantic coherence, and grounding performance across four benchmarks. All experiments use LLaVA-v1.5-13B as the base model and adopt greedy decoding with temperature set to 0 for consistency. Bold results indicate the best-performing configuration.
@@ Appendix B Algorithm Flowchart
Appendix C Extended Ablation Studies
C.1 Impact of token-level perturbation magnitude
We vary the token-level perturbation ratio $\omega$ and report results in Table 4 to investigate how perturbation strength affects model performance. In our method, Equation (9) governs the selection of tokens for perturbation based on their visual irrelevance. Specifically, we compute the similarity between visual features $\mathcal{G}_{v}(x)$ and text token embeddings $\mathcal{G}_{t}(q_{i})$ to estimate token-level visual alignment. Tokens with the lowest scores are considered visual-agnostic and thus are most eligible for perturbation. The perturbation budget $N_{t}$ is adaptively determined via the scaling coefficient $\omega$ and model uncertainty $\Delta P$ as:
$\mathcal{A}=\text{Top}_{N_{t}}(-\mathcal{G}_{v}(x)\cdot\mathcal{G}t(q_{i})^{T}),\quad N_{t}=\left\lfloor\omega\cdot\Delta P^{-1}\right\rfloor+1,$ (21)
where $\Delta P=\max_{j}P_{j}-\max{k\neq j}P_{k}$ quantifies the margin between the top two token-level alignment scores. This formulation encourages stronger perturbations under high uncertainty and milder changes when the model is confident.
As shown in Table 4, moderate values of $\omega$ lead to optimal hallucination suppression across both AMBER and OBJHal. Excessively low or high perturbation strengths either under-regularize or destabilize training. When $\omega$ is too small (e.g., $1\mathrm{e}{-4}$), the induced distributional shift is limited, resulting in marginal improvement over the baseline. This insufficient perturbation fails to adequately expose the model’s reliance on spurious token-level correlations, leading to suboptimal alignment correction. Conversely, overly large values (e.g., $5\mathrm{e}{-3}$ or $1\mathrm{e}{-2}$) introduce excessive perturbation into visual regions, disrupting the semantic coherence of inputs. This degrades both hallucination control and downstream task accuracy, as the model overfits to unstable signals. The best results are obtained at $\omega=1\mathrm{e}{-3}$, which achieves a balance between perturbation diversity and input integrity.
C.2 Impact of frequency regularization
To assess the contribution of spectral regularization, we conduct an ablation study by varying the frequency loss weight $\lambda$ in the TARS objective (Equation (12)). The term $\mathcal{L}_{\text{freq}}$ encourages semantic consistency between original and perturbed representations by aligning their frequency spectra. While conceptually appealing, the strength of this constraint is modulated by $\lambda$ and may interact non-trivially with preference supervision.
We evaluate $\lambda\in\{0.01,0.02,0.05,0.10,0.20,0.50,1.00\}$ and present results in Table 5. We omit $\lambda=0.00$ here as it reduces the method to DPO with perturbation. As shown in the table, introducing spectral alignment leads to consistent improvements across all benchmarks. Performance improves steadily as $\lambda$ increases from 0.01 to 0.20, with hallucination rates (CHAIR, Hal-Rate, $_{\text{s}}, CR$$_{\text{i}}$) decreasing and coverage improving. The best trade-off is achieved at $\lambda=0.20$, where TARS achieves the lowest hallucination and strongest grounding.
Beyond this point, we observe diminishing or adverse effects. For example, at $\lambda=0.50$ and $\lambda=1.00$, performance begins to degrade, particularly on MMHal and OBJHal. This trend suggests that overly aggressive regularization may constrain the model’s ability to accommodate subtle semantic variations introduced by token-level perturbations, leading to underfitting or conservative outputs.
These results confirm that spectral alignment is an effective regularizer when applied with moderate strength. It improves semantic coherence across perturbed samples without rigidly enforcing token-level correspondence, thus allowing preference optimization to remain robust yet expressive under controlled distributional shifts.
Table 5: Ablation study on the effect of spectral alignment weight $\lambda$. We evaluate the impact of varying $\lambda$ on hallucination suppression and multimodal alignment. All experiments are conducted using LLaVA-v1.5-13B as the base model and employ greedy decoding with temperature set to 0 for consistency. Bold numbers indicate the best across each metric.
Spectral Coefficient AMBER MMHal POPE OBJHal
@@ C.2 Impact of frequency regularization
Appendix D Additional Experimental Results

### v4

@@ 5 Related Work
Algorithm Data Size Feedback Reward-Free Hal. Mitigation Causal Align.
LLaVA-v1.5 [llava_origin] - -
+ RLHF [RLHF] 122k self-reward
+ RLAIF [RLAIF] 16k LLaVA-Next
+ HALVA [HALVA] 22k GPT-4V
+ DPO [li2024multi] 5k self-reward
+ CHiP-DPO [CHIP] 5k self-reward
+ OPA-DPO [OPA_DPO] 4.8k GPT-4V
@@
Multimodal large language models (MLLMs) extend LLMs by integrating visual inputs to support multimodal reasoning [chen2024mllm, reasoning, jain2024vcoder]. Typically, visual features are extracted by a vision encoder, aligned through a connector, and processed by the LLM [llava_origin, parekh2024concept]. Despite strong performance, MLLMs often produce factually incorrect or visually ungrounded outputs, undermining reliability [bai2024hallucination, chen2025perturbollava]. This issue is more severe than in unimodal LLMs [chen2024multi, jiang2024hallucination], mainly due to modality imbalance [ma2024visual, he2024ma] and ineffective fusion [bellagente2023multifusion, ji2023map]. Recent studies attribute these failures to persistent misalignment between multimodal representations and human expectations, rather than model capacity [chen2024unified, liu2024aligning, hao2025hccm].
@@
A key bottleneck in addressing MLLM hallucinations lies in aligning model outputs with human preferences for factual consistency. Unlike knowledge-intensive pretraining [chang2024large, mckinzie2024mm1] and instruction tuning [chen2024coin, llava_origin], recent methods typically leverage small-scale human preference data refined via reinforcement learning [yu2024rlhf, casperopen, feng2025rewardmap]. Direct preference optimization (DPO) [rafailov2023direct, DPO_MLLMs] has become a leading approach due to its simplicity and effectiveness, demonstrated in CHiP-DPO [CHIP] and OPA-DPO [OPA_DPO]. However, DPO’s reliance on limited data can cause overfitting to superficial linguistic cues [setlur2024rl, CHIP], leading to distributional rigidity and reduced adaptability to modality-specific semantics [ouali2024clip, song2024importance]. These limitations call for more adaptive alignment strategies that capture token-level variability and cross-modal dependencies for stable multimodal reasoning. Adversarial training has proven effective for improving robustness in vision-language pretraining [gan2020large, zhang2025towards], yet its integration into preference optimization for hallucination mitigation remains largely unexplored.
@@
To address these challenges, we propose a token-adaptive min-max alignment strategy with spectral regularization that enhances preference learning without relying on high-resource expert feedback (e.g., GPT-4V [achiam2023gpt]). Our spectral alignment loss operates in the frequency domain via FFT, encouraging global semantic consistency while tolerating the local distributional shifts induced by token perturbation. Using only a small public preference dataset, our method effectively mitigates hallucinations and consistently outperforms RL-based baselines across benchmarks. Tab.˜4 compares preference optimization methods in terms of data scale, supervision, and alignment.
@@ 6 Conclusion
This work reveals a fundamental limitation of static preference optimization: its tendency to latch onto superficial linguistic correlations rather than causally grounded visual evidence. We address this not by scaling data, but by reshaping the optimization landscape itself. TARS recasts preference learning as a min-max game in which the inner adversarial maximization over visual-agnostic tokens forces the model to disentangle causal visual signals from spurious textual cues, while a spectral alignment loss in the frequency domain provides a theoretically principled regularizer that preserves global semantic invariance under the resulting distributional shifts. Empirically, using only 4.8k preference samples without expert feedback, TARS halves the hallucination rate (26.4% $\to$ 13.2%), reduces cognitive inconsistency by an order of magnitude (2.5 $\to$ 0.4), and surpasses $5\times$ LLM-based augmentation on 28.8k samples while narrowing the gap with GPT-4o at a fraction of the model scale. These results establish that principled adversarial token perturbation can rival brute-force data scaling for robust multimodal preference alignment.
@@
  • Section 7 details model configurations, training settings for DPO and TARS, and token perturbation procedures.
  • Section 8 presents the min-max optimization algorithm of TARS in pseudocode form.
  • Section 9 includes extended ablation studies, covering perturbation magnitude, spectral regularization strength, a formal theoretical analysis of spectral alignment (energy dispersion and low-frequency masking bounds), spectral regularization alternatives, adaptive vs. random token selection, and min–max formulation analysis.
  • Section 10 reports additional benchmark results, including fine-grained AMBER hallucination metrics, extended MS-COCO CHAIR and multimodal understanding evaluations, comparison with data augmentation baselines, and comparison with retrieval-augmented DPO.
  • Section 11 presents results on Muffin-13B, demonstrating the generality of TARS across different MLLM architectures.
  • Section 12 discusses model behavior, including sensitivity and design insights.
  • Section 13 showcases qualitative comparisons on representative examples.
7 Implementation Details
7.1 Base Model Setups
We evaluate our method on LLaVA-v1.5 [llava_origin] models with 7B and 13B parameters. LLaVA-v1.5 adopts Vicuna-7B/13B [chiang2023vicuna] as the language backbone and CLIP-ViT-L/14 [radford2021learning] as the vision encoder. The vision encoder also serves as the similarity function $\mathcal{G}(\cdot)$ used in Eq.˜9 to compute alignment between visual inputs and text tokens. All experiments are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
7.2 DPO Training Setups
@@
7.3 Evaluation Benchmark Setups
We evaluate TARS on four established hallucination benchmarks:
1) AMBER [wang2023amber] (Generative): A fine-grained benchmark for hallucination evaluation. Following prior works [OPA_DPO, CHIP], we evaluate only the generative subset using the official codebase. Metrics include CHAIR [rohrbach2018object] for object hallucination detection, object coverage (Cover) for completeness, response-level hallucination rate (Hal-Rate), and alignment with human cognition (Cog).
2) MMHal [sun2023aligning] (Generative): A VQA benchmark with real-world scenarios, evaluated using GPT-4V feedback to measure overall scores and hallucination rates (Hal-Rate).
3) OBJHal [yu2024rlhf] (Generative): A benchmark evaluating hallucinations in image captioning. We report hallucination rates at the response level (CRs) and object mention level (CRi).
4) POPE [li2023evaluating] (Discriminative): A binary VQA benchmark designed to assess object hallucination through yes/no questions, using popular, random, and adversarial sampling strategies.
We follow the original evaluation settings and benchmark splits for AMBER, MMHal, and OBJHal as specified in their respective papers. For POPE, we construct a benchmark of 9,000 VQA pairs by sampling using the three strategies above.
For evaluation metrics, we adopt four response-level hallucination measures across different benchmarks: CHAIR [rohrbach2018object] for object hallucination detection, object coverage (Cover) for completeness measurement, response-level hallucination rate (Hal-Rate) for overall hallucination assessment, sentence-level hallucination rate (CRs) for holistic response evaluation, and object mention-level hallucination rate (CRi) for fine-grained object-level analysis.
For evaluation feedback collection, we employ the en-core-web-lg English NLP pipeline for AMBER to extract structured semantic cues as lightweight and reproducible evaluators. For MMHal and OBJHal, we utilize the expert GPT-4V model [hurst2024gpt] (gpt-4-1106-vision-preview) for feedback evaluation, following the established protocols.
8 Algorithm Flowchart
@@
4:     Compute token-level visual relevance: $P_{i}=\mathcal{G}_{v}(x)\cdot\mathcal{G}_{t}(q_{i})^{T}$.
5:     Estimate model confidence margin: $\Delta P=\max_{j}P_{j}-\max_{k\neq j}P_{k}$.
6:     Determine adaptive perturbation budget: $N_{t}=\left\lfloor\omega\cdot\Delta P^{-1}\right\rfloor+1$.
7:     Select visually agnostic tokens: $\mathcal{A}=\text{Top}_{N_{t}}(-P)$.
8:     Apply controlled perturbation to obtain $\varphi(q)$.
@@
10:     Compute the preference alignment loss $\mathcal{L}_{\text{DPO}}$ via DPO.
11:     Apply frequency-domain regularization $\mathcal{L}_{\text{freq}}$.
12:     Compute final objective: $\mathcal{L}_{\text{TARS}}=\mathcal{L}_{\text{DPO}}+\lambda\cdot\mathcal{L}_{\text{freq}}$.
@@
9 Extended Ablation Studies
9.1 Impact of Token-Level Perturbation Magnitude
We vary the token-level perturbation ratio $\omega$ and report results in Tab.˜5 to investigate how perturbation strength affects model performance. In our method, Eq.˜9 governs the selection of tokens for perturbation based on their visual irrelevance. Specifically, we compute the similarity between visual features $\mathcal{G}_{v}(x)$ and text token embeddings $\mathcal{G}_{t}(q_{i})$ to estimate token-level visual alignment. Tokens with the lowest scores are considered visual-agnostic and thus are most eligible for perturbation.
As shown in Tab.˜5, moderate values of $\omega$ lead to optimal hallucination suppression across both AMBER and OBJHal. Excessively low or high perturbation strengths either under-regularize or destabilize training. When $\omega$ is too small (e.g., $1\mathrm{e}{-4}$), the induced distributional shift is limited, resulting in marginal improvement over the baseline. Conversely, overly large values (e.g., $5\mathrm{e}{-3}$ or $1\mathrm{e}{-2}$) introduce excessive perturbation into visual regions, disrupting the semantic coherence of inputs. The best results are obtained at $\omega=1\mathrm{e}{-3}$, which achieves a balance between perturbation diversity and input integrity.
Table 5: Ablation study on the effect of token-level perturbation magnitude $\omega$ in TARS. All experiments use LLaVA-v1.5-13B as the base model. Bold results indicate the best-performing configuration.
@@ 9.1 Impact of Token-Level Perturbation Magnitude
9.2 Impact of Frequency Regularization
To assess the contribution of spectral regularization, we conduct an ablation study by varying the frequency loss weight $\lambda$ in the TARS objective (Eq.˜12). We evaluate $\lambda\in\{0.01,0.02,0.05,0.10,0.20,0.50,1.00\}$ and present results in Tab.˜6.
Performance improves steadily as $\lambda$ increases from 0.01 to 0.20, with hallucination rates decreasing and coverage improving. The best trade-off is achieved at $\lambda=0.20$. Beyond this point, performance begins to degrade, particularly on MMHal and OBJHal. This trend suggests that overly aggressive regularization may constrain the model’s ability to accommodate subtle semantic variations introduced by token-level perturbations.
Table 6: Ablation study on the effect of spectral alignment weight $\lambda$. All experiments are conducted using LLaVA-v1.5-13B. Bold numbers indicate the best across each metric.
Spectral Coeff. AMBER MMHal POPE OBJHal
@@ 9.2 Impact of Frequency Regularization
9.3 Theoretical Analysis of Spectral Alignment
We provide a formal justification for adopting frequency-domain alignment instead of spatial losses under token-level adversarial perturbations in TARS.
Setup. Let a hidden-state sequence be $\mathbf{Z}=(\mathbf{z}_{0},\dots,\mathbf{z}_{L-1})\in\mathbb{R}^{L\times d}$. A single-token adversarial perturbation at position $m$ induces a corrupted sequence $\tilde{\mathbf{Z}}$:
$\tilde{\mathbf{z}}_{i}=\mathbf{z}_{i}+\boldsymbol{\delta}\cdot\mathbb{I}(i=m),$ (13)
where $\boldsymbol{\delta}\in\mathbb{R}^{d}$ is bounded ($\|\boldsymbol{\delta}\|\leq\epsilon$) and $\mathbb{I}(\cdot)$ is the indicator function. The Discrete Fourier Transform (DFT) along the token axis maps $\mathbf{Z}$ into the frequency domain:
$\hat{\mathbf{z}}_{k}=\sum_{i=0}^{L-1}\mathbf{z}_{i}\,e^{-j\frac{2\pi}{L}ki},\quad k=0,\dots,L{-}1,$ (14)
where $j$ denotes the imaginary unit.
Proposition 1 (Spectral Energy Dispersion of Local Perturbation). Under a single-token perturbation at position $m$, the frequency-domain deviation is:
$\tilde{\hat{\mathbf{z}}}_{k}=\hat{\mathbf{z}}_{k}+\boldsymbol{\delta}\,e^{-j\frac{2\pi}{L}km}.$ (15)
Taking the $\ell_{2}$ norm yields:
$\|\tilde{\hat{\mathbf{z}}}_{k}-\hat{\mathbf{z}}_{k}\|_{2}=\|\boldsymbol{\delta}\|_{2}\quad\forall\,k\in\{0,\dots,L{-}1\}.$ (16)
Proof. By linearity of the DFT, $\tilde{\hat{\mathbf{z}}}_{k}-\hat{\mathbf{z}}_{k}=\boldsymbol{\delta}\,e^{-j\frac{2\pi}{L}km}$. Since the complex exponential has unit modulus ($|e^{-j\theta}|=1$), the $\ell_{2}$ norm of a real vector scaled by it is unchanged. $\blacksquare$
Remark. Proposition 1 reveals that a concentrated spatial spike (single-token perturbation) acts as a Dirac-like impulse whose energy is uniformly dispersed across all frequency bins $k$.
Theorem 1 (Bounded Semantic Shift via Low-Frequency Masking). Let the spatial alignment loss be $\mathcal{L}_{\mathrm{spatial}}=\|\tilde{\mathbf{Z}}-\mathbf{Z}^{\mathrm{ref}}\|_{F}^{2}$. By Parseval’s theorem,
$\mathcal{L}_{\mathrm{spatial}}=\frac{1}{L}\sum_{k=0}^{L-1}\|\tilde{\hat{\mathbf{z}}}_{k}-\hat{\mathbf{z}}^{\mathrm{ref}}_{k}\|_{2}^{2}.$ (17)
To decouple global semantics from local token noise, TARS introduces a low-frequency alignment loss with a binary mask $\mathcal{M}$ ($\mathcal{M}_{k}=1$ for $k<K_{\mathrm{cut}}$, and 0 otherwise):
$\mathcal{L}_{\mathrm{freq}}=\frac{1}{L}\sum_{k=0}^{L-1}\mathcal{M}_{k}\,\|\tilde{\hat{\mathbf{z}}}_{k}-\hat{\mathbf{z}}^{\mathrm{ref}}_{k}\|_{2}^{2}.$ (18)
Under the perturbation $\boldsymbol{\delta}$, the frequency loss is upper-bounded by:
$\mathcal{L}_{\mathrm{freq}}\leq\frac{K_{\mathrm{cut}}}{L}\,\|\boldsymbol{\delta}\|^{2}<\mathcal{L}_{\mathrm{spatial}}.$ (19)
Proof. Applying Proposition 1, each masked bin contributes at most $\|\boldsymbol{\delta}\|^{2}/L$ to $\mathcal{L}_{\mathrm{freq}}$. Summing over $K_{\mathrm{cut}}<L$ bins gives the bound. $\blacksquare$
Interpretation for Min-Max Optimization. In the inner maximization of TARS, the model generates local adversarial shifts $\boldsymbol{\delta}$ to explore hallucination boundaries.
  • Limitation of spatial loss. Enforcing spatial alignment penalizes $\|\boldsymbol{\delta}\|^{2}$ entirely at the perturbed position $m$. This rigid positional constraint restricts the adversarial exploration space that the inner maximization is designed to exploit.
  • Advantage of spectral loss. Because the dominant semantic structure of MLLM hidden states concentrates in the low-frequency modes (i.e., $\|\hat{\mathbf{z}}_{k}\|$ is substantially larger for small $k$), the relative distortion $\|\boldsymbol{\delta}\|/\|\hat{\mathbf{z}}_{k}\|$ in the masked low-frequency regime approaches zero. Thus, $\mathcal{L}_{\mathrm{freq}}$ acts as a robust invariant: it forces the global semantic envelope (low frequencies) to align with the non-hallucinated reference, while safely ignoring the high-frequency positional noise induced by the min-max perturbation.
9.4 Analysis of Spectral Regularization Alternatives
FFT-based regularization decomposes hidden representations into low- and high-frequency components along the token dimension. (1) Low-frequency alignment preserves global semantic structure (e.g., topic, intent, and visual grounding) under token-level perturbations, since the dominant low-frequency modes capture the overall distributional statistics of the hidden-state sequence. (2) High-frequency constraint mitigates overfitting to local lexical patterns and positional artifacts, improving robustness and generalization. (3) Why FFT over spatial metrics? Given hidden states $H=[h_{1},\dots,h_{T}]\in\mathbb{R}^{T\times D}$, an $\ell_{2}$ loss $\|H-H^{\prime}\|_{2}$ treats every positional deviation equally, amplifying position-dependent noise from our token perturbations. In contrast, the FFT representation $\mathcal{F}(H)$ aggregates information across all positions into frequency bins; a local perturbation at position $l$ contributes a bounded additive term $\propto e^{-2\pi ikl/L}$ to each frequency bin $k$, resulting in smooth spectral variation. As shown in Tab.˜7, FFT-based regularization consistently outperforms $\ell_{2}$ and cosine baselines across all metrics, validating this design choice. Additionally, Tab.˜15 compares spectrum-based alignment (TARS) with token-level contrastive alignment under identical perturbation policies, showing that spectral alignment achieves substantially lower hallucination rates (13.2% vs. 16.3%) and higher coverage (59.6% vs. 57.2%).
Table 7: Ablation on spectral regularization alternatives. FFT-based regularization outperforms $\ell_{2}$ and cosine baselines.
[…]
$\emptyset$ 2.8 58.3 15.1 0.7 2.29 0.52
$\ell_{2}$ 2.7 58.5 15.4 0.9 2.32 0.49
Cosine 2.5 58.8 14.5 0.7 2.34 0.49
FFT 2.4 59.6 13.2 0.4 2.48 0.45
9.5 Adaptive vs. Random Token Selection
A critical design choice in TARS is the adaptive selection of perturbation targets based on cross-modal alignment scores, rather than random or uniform token selection. We validate this by comparing three strategies: (1) vision-only perturbation (perturbing tokens with high visual relevance); (2) random perturbation (uniformly sampling tokens to perturb); and (3) our proposed visual-agnostic perturbation (selectively perturbing tokens with the lowest cross-modal alignment).
[…]
Table 8: Comparison of perturbation target strategies. Visual-agnostic perturbation (TARS) consistently outperforms alternatives.
[…]
LLaVA-v1.5-7B 7.6 51.7 35.4 4.2 2.02 0.61
+ DPO 4.9 56.6 26.4 2.5 2.19 0.61
+ Vision-only Perturb. 5.8 55.2 31.6 3.8 2.12 0.66
+ Random Perturb. 4.4 57.8 28.7 3.1 2.30 0.64
+ TARS (ours) 2.4 59.6 13.2 0.4 2.48 0.45
9.6 Min–Max Formulation Analysis
A key design choice in TARS is the adversarial min–max formulation (Eq.˜6), which replaces the static preference optimization of standard DPO with a dynamic two-stage procedure. The full empirical analysis of training dynamics is presented in the main text (Sec.˜4.6, Fig.˜7). Here we provide additional protocol details.
We train all three variants (standard DPO, DPO with random perturbation (RandPert), and TARS (Min-Max)) on LLaVA-v1.5-7B with identical hyperparameters using 4.8k preference samples over 20 epochs. All reported loss values are normalized to $[0,1]$ to enable fair cross-method comparison. The test set is evaluated under two conditions: clean (original held-out samples) and robust (token-perturbed held-out samples), isolating generalization from robustness. The results confirm that: (1) DPO achieves the lowest training loss ($\sim$0.2) but the worst robust test loss (0.65–0.75), exhibiting severe overfitting; (2) RandPert yields intermediate but highly unstable convergence; and (3) the Min-Max formulation achieves the best test performance on both clean (<<0.3) and robust (<<0.5) splits, acting as an implicit regularizer that prevents premature convergence to spurious patterns.
10 Additional Experimental Results
10.1 AMBER Generative and Discriminative Metrics